OpenStudy (anonymous):

In a class of 120 students numbered 1 to 120, all even number students opt for physics, whose number are divisible by 5 opt for chemistry & those whose no are divisible by 7 opt for math. How many opt for none of these subjects?

OpenStudy (anonymous):

79 do not opt for these subjects

OpenStudy (anonymous):

u hv 24 multiples of 5 from 1 to 120 n 17 multiples of 7 so 120 - (24+17) = 79

OpenStudy (anonymous):

what about even numbered students isnt there 60 of these?

OpenStudy (anonymous):

there are 60 even numbers ...hence 60 physics student.. then 12..opt for chemistry...as 5 ,15,25,.....115(we cann't select 10,20..as it is even and seleccted in physics) we have 17*7=119 ..i.e 17 but we cann't take 35,70,105...because they are selected earlier hence rest are=120-(60+12+14)=120-86=34

OpenStudy (anonymous):

we have 17 ...7 but as 35,70,105 is not ther hence 14...deducted

OpenStudy (anonymous):

lol - this is harder than it seems

OpenStudy (anonymous):

yep prashant - u r right

OpenStudy (anonymous):

this question was hard...

OpenStudy (anonymous):

nobody give right answer someone reach at last but not find correct answer

OpenStudy (anonymous):

ohh god!!

OpenStudy (anonymous):

i hope any one give true answer

OpenStudy (anonymous):

Sorry I misread the question in a hurry Is 41 the right answer???

OpenStudy (anonymous):

u r right prashant tell me explenation

OpenStudy (anonymous):

there r 60 even numbers between 1 n 120 there r 24 numbers divisible by 5 between 1 n 120. however 12 of them end in 0 so r even numbers and so ruled out. so multiples of 5 left r 12. there r 17 multiples of 5 between 1 n 120. however 8 of them r multiples of 2 and 2 r multiples of 5. so multiples of 7 left r 7. total works out to 60 + 12 + 7 = 79 so left r 120 - 79 = 41

OpenStudy (anonymous):

ya harkirat is right....although my approach is right...but i did mistake in case of 7.... there are some even like 28,42...are removed....

OpenStudy (anonymous):

good solution master

OpenStudy (anonymous):

U r welcome, i was in great the first time around n misread.....☺

OpenStudy (anonymous):

ok genious

OpenStudy (anonymous):

prashant your approch also right but at the end u doing mistake

OpenStudy (anonymous):

harkirat i expect from u that u solve my unanswered question...........its level is too much high........try it

OpenStudy (anonymous):

the one about factors ????

OpenStudy (anonymous):

yes it is to much typical? and also there one of mangoes problem